Circular No. 5733

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993K IN NGC 2223 A. Williams, University of Western Australia; and R. Martin, Perth

Observatory, report their discovery of a possible supernova of mag 14.5 ∼ 30′′ east and 31′′ north of the nucleus of NGC 2223 (α = 6^{h}22^{m}.5, δ = −22^{o}49′, equinox 1950.0). The discovery was made on Mar. 28.55 UT in the course of the Perth Observatory Research Group’s automated supernova search program, using the 0.61-m Perth-Lowell automated telescope. The object was reobserved on Mar. 29.5, but it had not been present on Mar. 10.5 UT.

P. C. Schmidtke, Arizona State University, obtained BV RI images of the field with the 0.9-m telescope at the Cerro Tololo Interamerican Observatory, starting on Mar. 30.09 UT. He confirmed the presence of a star ∼ 30′′ east and 28′′ north of the center of the galaxy, quick-look photometry yielding V = 15.4, B − V = +0.2, V − R = +0.1.

M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that J. Maza and M. Wischnjewsky, University of Chile, obtained a low-resolution spectrum (range 320-760 nm) of the object with the CTIO 1.5-m telescope. The spectrum is essentially featureless, suggesting that this supernova—like 1993J—is a type II event in an early stage.

SUPERNOVA 1993J IN NGC 3031 G. Gomez, Instituto de Astrofisica de Canarias; and R. Lopez, Univer-

sity of Barcelona, report: “A spectrum (range 500-850 nm, resolution 0.3 nm) taken on Mar. 30 by E. Perez with the 2.5-m Isaac Newton Telescope shows some similarity to the premaximum spectrum of the type II supernova 1990E. In particular, there is a very broad (FWZI ∼ 70 nm) emission band toward 650 nm that could correspond to Hα.”

Corrigenda. On IAUC 5731, the names P. Pujol and B. A. Hong should have been given as F. Pujol and H. B. Ann, respectively, and the time of the first CCD image should have been Mar. 28.89, rather than Mar. 29.1.

Further visual magnitude estimates: Mar. 29.95 UT, 11.1 (F. Garcia, Lugo, Spain); 30.84, 10.3 (P. Schmeer, Bischmisheim, Germany).

1993 March 30 Brian G. Marsden